OpenStudy (natalie1234):

PLEASE HELP! WILL GIVE MEDAL Fill in the missing information in the problem to find square root of 175 by using the Babylonian method. Let 10.5 be your initial guess.

OpenStudy (anonymous):

that is a really really lousy first guess

OpenStudy (anonymous):

\[10^2=000\\ 11^2=121\\ 12^2=144\\ 13^2=169\]

OpenStudy (anonymous):

you want \(\sqrt{175}\) start with \(13\) it is much much closer than \(10.5\) a very dumb far away guess

OpenStudy (anonymous):

wanna do it with 13?

OpenStudy (natalie1234):

okay haha well they used 10.5 in the question

OpenStudy (natalie1234):

sure

OpenStudy (anonymous):

idiots ok well with the first guess of \(10.5\) the second guess will be \[\frac{1}{2}\left(10.5+\frac{175}{10.5}\right)\]

OpenStudy (anonymous):

if the first guess is 13 the second guess will be \[\frac{1}{2}\left(13+\frac{175}{13}\right)\]

OpenStudy (anonymous):

either way you probably want to use a calculator

OpenStudy (natalie1234):

ok

OpenStudy (anonymous):

course i am not sure what the "missing information" is you are supposed to fill in

OpenStudy (anonymous):

ok i think there was a typo in your question i bet it said find \(\sqrt{170}\) with an initial guess of \(13.5\) which makes a lot more sense

OpenStudy (natalie1234):

oh well thats how they gave me the question just like i wrote it

OpenStudy (natalie1234):

but i think they did do a typo lol wow

OpenStudy (anonymous):

the set up is for the question i wrote the answer is \[\frac{13.5+\frac{170}{13.5}}{2}\]

OpenStudy (anonymous):

i.e. put a \(13.5\) where you see the question mark

OpenStudy (natalie1234):

ok